UTC 2011 May 17

Observer(s): T. Sargent, P. Grenz, A. Rakich, J. Hill (LBTO, Tucson)
Telescope Operator: D. Huerta (LBTO)
Telescope Support: J. Little (LBTO)
Instrument Support: J. Morris (LBTO)

Summary

Ilya got monocular presets and offsets sorted out from the PEPSI interface.

We lost half of the first half to wind, but still got a decent amount of on-sky testing for BP11.

We took a few frames of LUCI1@LFBG and IRTC2@RFBG to test the Sandbox.binorb.cl script.

We sorted out some BP11 problems for LBC just in time to be closed by wind for the third time.

Preparations

Installed new version of BP11 to fix some issues from last night.

Details

TCS Testing with BP11 (first half)

Lost first half hour to wind.

Ilya got monocular presets and offsets working from his interface on both sides of the telescope.

We made significant progress on the monocular on-sky portions of TCS testing.

Closed for wind at 04:10 UT.

LBC software not working with BP11 (lbcserver crashes after IIF_Authorize completes.)
Paul fixed this (temporary) by relinking the LBC code to a modified version of BP10.
A permanent fix was installed the next day.

Ran MODS presets using LBTtools and talking to MODS1 AGW.

Found that TCSSendWavefront wasn't working. Update the link to the new /lbt/tcs/....... path, but still issues. Paul and Stephen figured out that the variable iifIP = "iif" was missing from the new tcs.conf file. We did: cat iif.conf >> tcs.conf and that got things working. See Issue 3373.

Bent Gregorian Range Balancing Tests (second half)

Re-Open

08:20 Sync ACQUIRE Presets for LUCI1/IRTC2 (star was in the acquisition field on both sides with default pointing models)

08:22 Sync ACTIVE Presets for LUCI1/IRTC2

08:47 Async TRACK Presets (to measure the co-pointing and range balance)

irtc.20110517.085305.fits
luci.20110517.0002.fits
irtc.20110517.085921.fits
luci.20110517.0003.fits

Working on these images with Sandbox.binorb.cl in /scratch/20110517 on obs3. binorb runs, but fails when it is reading the rotator angles from the headers. I think it is failing because rotmode is 'unkown' in the luci headers. This may be because I sent the preset from IRAF rather than luci (no Telescope Commit in luci). (It was a preset in POSITION mode with PA=0)

LBC Tests (almost)

09:15 Switching to LBC

Close

09:35 Closed again for high winds just as we finish the switch.

Finish

Put the blinds on LUCI1 and handed it over to Walter for testing.

Turned off the MODS AGW (leaving the software running as I found it).

Stopped the IRTC cooling.

Turned off LBC cameras.

10:15 Gone for the night as we are closed for high winds again. Allowing time for Walter testing LUCI1, and Ilya sending some more presets.

-- JohnHill - 16 May 2011
Topic revision: r4 - 18 May 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback